OpenStudy (anonymous):

Eliminate the parameter to find the Cartesian Equation of the following curve? \(\ \Huge x=ln(t), y=\sqrt{t}, t \ge 1\)

OpenStudy (experimentx):

x = ln(y^2)

OpenStudy (anonymous):

How do I do this problem, though? I have a little difficulty with \(\ \Huge e \) and \(\ \Huge ln. \)

OpenStudy (experimentx):

\[ \huge x = 2 \ln y \\ \huge y = e^{x \over 2} \]

OpenStudy (anonymous):

@experimentX Where does \(\ \Huge y=e^{\frac{x}{2}} \) come from \(\ \Huge ? \)

OpenStudy (experimentx):

 e and ln are inverse of each other.